OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (-5, 1) with a slope of three halves. y - 5 = three halves(x + 1) y + 1 = three halves(x - 5) y + 5 = -three halves(x - 1) y - 1 = three halves(x + 5)

OpenStudy (tamtoan):

with slope of 3/2 ...use this formula : m = \[\frac{ y - y1 }{ x - x1 } with point (x1,y1) = (-5,1) \] to see if you can work out the equation ? m is the slope